Mulbiple Production Department Factory Overhead Rates Spotted Cow Dairy Company manufectures three products-whole milk, skdm milk, and cream-in two production departments, Biending and Packing. The factory overhead for Spotted Cow Dairy is $340,100 The three products censume both machine hours and direct labor hours in the two preduction departments as follows Direct Labor Hours Machine Hours 610 390 300 1,300 830 Sim mik 530 2,030 Packing Department Whole mik 1,280 190 175 545 1,845 780 2,900 4,930 The management of Spotted Cow Dairy Company now plans to use the multiple production department factory overhead rate method. The total factory overhead associated with each department is as follows: Biending Department Packing Department $ 1 69,000 171,100
media%2F8d4%2F8d4fdc2b-0a62-4191-8b59-9e
Answer #1

Formula sheet

A B C D E F G H I J K L
2
3 Calculation of overhead rate for Blending department:
4 Allocation Base Direct Labor Hours
5 Total Overhead for Blending Department 169000
6 Total direct labor hours for blending department 1300
7
8 Overhead Rate for Blending department =Total Overhead for blending deaprtment / Total Direct Labor hours for Blending Department
9 =D5/D6 =D5/D6
10
11 Hence Overhead Rate for Blending department =D9
12
13
14 Calculation of overhead rate for Packing department:
15 Allocation Base Machine Hours
16 Total Overhead for Packing Department 171100
17 Total Machine hours for packing department 2900
18
19 Overhead Rate for Packing department =Total Overhead for Packing deaprtment / Total Machine hours for Packing Department
20 =D16/D17 =D16/D17
21
22 Hence Overhead Rate for Packing department =D20
23
24
25 Calculation of product factory overhead costs:
26 Whole Milk Skim milk Cream
27 Number of labor hours used in Blending Department 610 390 300
28 Overhead Rate for Blending Department =$D$11 =$D$11 =$D$11
29 Blending Department factory overhead =D27*D28 =E27*E28 =F27*F28
30
31 Whole Milk Skim milk Cream
32 Number of machine hours used in Packing Department 1280 780 840
33 Overhead Rate for Packing Department =$D$22 =$D$22 =$D$22
34 Packing Department factory overhead =D32*D33 =E32*E33 =F32*F33
35
36 Hence,
37 Whole Milk Skim milk Cream
38 Blending Department factory overhead =D29 =E29 =F29
39 Packing Department factory overhead =D34 =E34 =F34
40 Total factory Overhead =D38+D39 =E38+E39 =F38+F39
41